What is the surface area of a cube that has an edge measuring 5 inches?
Found 3 solutions by edjones, dabanfield, richwmiller:
Answer by edjones(8007)

5*5=25 area of one side
25*6=150 in^2 surface area (6 sides)
.
Ed

Answer by dabanfield(803)

What is the surface area of a cube that has an edge measuring 5 inches?
A cube has six faces each of which is a square. In this case each side of each square is 5 inches so each face has surface area = 5*5 = 25 square inches.
Since there are six faces we have a total surface area = 6*25 = 150 square inches.

Answer by richwmiller(17219)

all the sides of a cube are the same.
so surface area will be 6 times the area of one side
area of a side is the area of a square of 5 inches on each side